For a mean of 153 and a standard deviation of 25, find P(130 ≤ x ≤ 170 )
1
Expert's answer
2020-06-01T18:00:23-0400

When the distribution is not mentioned, it is conventionally assumed to be Normal.

Here the distribution of X has not been mentioned, so we assume

X ~ N(μ = 153, σ2 = 252)


Then by the property of Normal distribution,


Z = "\frac{x-\mu}{\sigma}"​ ~ N(0, 1) i.e. Z is a standard normal variable


Therefore, P(130 ≤ x ≤ 170)


= P("\frac{130-153}{25}\leq\frac{x-153}{25}\leq\frac{170-153}{25}")


= P(- 0.92 ≤ Z ≤ 0.68)


= P(Z ≤ 0.68) - P(Z < - 0.92)


= "\Phi"(0.68) - "\Phi"(- 0.92)


= 0.7517 - 0.1788 = 0.5729


Answer: The value of P(130 ≤ x ≤ 170) is 0.5729.